use substitution to solve the liner system y=2x-2 and 5x-2y=18
Answer by rfer(16322)

5x-2(2x-2)=18
5x-4x+4=18
x=14
----------------
y=2(14)-2
y=26
-----------------
5*14-2*26=18
70-52=18